OpenStudy (anonymous):

soft drink bottling plant can produce 120,000 cans of soft drinks a day. At this rate, how long will it take to produce 15,000 cans?

OpenStudy (anonymous):

helo??

OpenStudy (anonymous):

I mean help

OpenStudy (anonymous):

what time unit are you using, days, hour, min, or seconds

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Its just a math question my teacher made I have no idea what time or day she was thinkin of

OpenStudy (anonymous):

well its a rate so just simplify 15/120 = 1/8 which is 1/8 of a day if its per hour then 24/8 = 3 hr